write the equation of the line that passes through the points (2,1,3) and (1,4,-3). write all three forms of the equations of the line.


1
Expert's answer
2021-11-17T10:05:49-0500

There is three major forms of linear equations:

  • vector form: "(x,y,z)=(x{\scriptscriptstyle 0},y{\scriptscriptstyle 0},z{\scriptscriptstyle 0})+t(a,b,c)"
  • parametric form: "x=x{\scriptscriptstyle 0}+ta, y = y{\scriptscriptstyle 0}+tb,z=z{\scriptscriptstyle 0}+tc"
  • symmetric form: "{\frac {x-x{\scriptscriptstyle 0}} a}={\frac {y-y{\scriptscriptstyle 0}} b}={\frac {z-z{\scriptscriptstyle 0}} c}"

Where "x{\scriptscriptstyle 0}, y{\scriptscriptstyle 0},z{\scriptscriptstyle 0}" - point on that line

(a, b, c) - directing vector of a line

t is a parameter, "t \isin R"

(a, b, c) = (1 - 2, 4 - 1, -3 -3) = (-1, 3, -6)

Then

  • vector form: "(x,y,z)=(2,1,3)+t(-1,3,-6)"
  • parametric form: "x=2-t, y = 1+3t,z=3-6t"
  • symmetric form: "{\frac {x-2} {-1}}={\frac {y-1} 3}={\frac {z-3} {-6}}"